OpenStudy (anonymous):

Winnie rented a room at $250 per week. She also had to pay a deposit of $1200. Part A: Write a function to show the total amount of money f(n) that Winnie has to pay for renting the room for n weeks. (4 points) Part B: What is the total rent that Winnie has to pay for using the room for 12 weeks? (4 points) Part C: If the deposit amount decreases by $100, what would be the total rent for 6 weeks? (2 points) @Venny

OpenStudy (anonymous):

Ok so if it is 250 dollars a week and n is the amount of weeks it will be 250n.

OpenStudy (anonymous):

y=250n+1200 n=12 gives y=250(12)+1200=4200 New function of f(n)=250(n)+1100, along 6 weeks or n=6, give f(6)=250(6)+1100=2600

OpenStudy (anonymous):

1200 + $250 a week for n weeks that is simply f(n) = 1200 + 250n

OpenStudy (anonymous):

A deposit is added on so it will be +1200

OpenStudy (anonymous):

Yes now for the next part you plug 12 in for n

OpenStudy (anonymous):

3,000.

OpenStudy (anonymous):

You forgot to add the 1200 on to it

OpenStudy (anonymous):

4,200 why would i add that though?

OpenStudy (anonymous):

Because that is part of the cost when she originally rented the room.

OpenStudy (anonymous):

for C f(6)=250(6)+1100=2600 ?

OpenStudy (anonymous):

Correct good job!

OpenStudy (anonymous):

I have 1 more i will tag you a different question.